OpenStudy (anonymous):

Suppose a rectangle has length 4 and width 5. What is its perimeter? Suppose a rectangle has length 4 and width 5. What is its perimeter? @Mathematics

OpenStudy (diyadiya):

2(4+5)=2*9=18

OpenStudy (diyadiya):

perimeter of rectangle=2*length+2*width =2(length+width)

OpenStudy (lalaly):

agree^^

OpenStudy (diyadiya):

:))

OpenStudy (anonymous):

Sorry I don't quite understand, where does the two come from?

OpenStudy (diyadiya):

is this understandable ?